# mode=buildmissing

### v1

@@ 1 Introduction
Since Feynman’s 1982 vision of quantum computation feynman2018simulating, significant progress has demonstrated that quantum computers can leverage quantum mechanical principles to achieve fundamental computational advantages over classical methods (steane1998quantum; ladd2010quantum; preskill2012quantum; demarti2024decoding). Landmark quantum algorithms have demonstrated computational advantages, including exponential speedup for factoring (shor1994algorithms) and quadratic search improvement (grover1996fast). Recent experimental demonstrations of quantum supremacy have further validated quantum computing’s potential across diverse domains (arute2019quantum; zhong2020quantum; wu2021strong; huang2022quantum; madsen2022quantum; bao2023very; bluvstein2024logical; aghaee2025scaling). These advances promise to revolutionize cryptography (ekert1991quantum; bennett2014quantum), optimization (kadowaki1998quantum; bharti2022noisy), materials science (lloyd1996universal), and machine learning (huang2022quantum; cerezo2022challenges).
@@
Yet, for practical quantum computation to become a reality, errors on the physical level must be corrected with high confidence. Despite recent advances, quantum noise remains a major obstacle to practical quantum computing (google2023suppressing). These errors arise through numerous mechanisms: quantum gates cause unwanted errors due to imprecise implementation (fowler2012surface), while additional errors stem from imperfections in the equipment (preskill2018quantum), interaction with the surrounding environment (burnett2019decoherence; etxezarreta2021time), or measuring quantum systems (fowler2012surface). While fault-tolerant quantum computation can theoretically be achieved through redundancy by combining multiple physical qubits into one logical qubit (shor1995scheme; nielsen2010quantum), this approach creates a critical computational bottleneck: QEC requires real-time decoding algorithms that must process syndrome measurements and determine corrections within microsecond timescales while maintaining near-optimal accuracy (terhal2015quantum; higgott2022pymatching). Current decoding methods face a fundamental trade-off between computational efficiency and error-correction performance, methods like MWPM (fowler2013minimum), Belief Propagation with Ordered Statistics Decoding (BP-OSD) (roffe2020decoding) and tensor network decoder (bravyi2014efficient) scaling prohibitively with code distance while faster heuristics sacrifice the accuracy essential for fault-tolerant operation (demarti2024decoding). The field of QEC has advanced significantly, with several families of QEC codes proposed, including topological codes (kitaev2003fault; bombin2006topological; fowler2012surface; chamberland2020topological), Quantum Low-Density Parity Check (QLDPC) codes (mackay2004sparse; panteleev2021quantum; breuckmann2021quantum), and quantum turbo codes (poulin2009quantum). Recently, there has been significant growth in machine learning techniques applied to quantum decoding (wang2024artificial; klusch2024quantum). However, existing neural decoders typically fail to achieve near-optimal error thresholds, creating a gap between the theoretical potential of learned approaches and the performance requirements of fault-tolerant quantum computing. We address this challenge by introducing a unified framework that combines transformer-based neural decoding with specialized architectural innovations. Our approach leverages neural networks to learn syndrome-to-error map- pings while employing dual-stream processing and logical-centric loss design to directly optimize logical error suppression. To achieve this, our framework introduces several key innovations:
@@
  • A novel dual-stream transformer architecture (vaswani2017attention) that simultaneously processes syndrome and logical information streams with specialized attention mechanisms, featuring global tokens (zaheer2020big) and structured masking patterns that capture the geometric constraints and local correlations inherent in stabilizer codes.
@@
Our results significantly outperform existing neural decoders like QEC Transformer (QECCT) (choukroun2024deep) and classical methods like MWPM, while matching the performance of computationally expensive approaches across both toric and rotated surface codes.
@@ 2 Related works
A broad suite of QEC codes has been devised to protect quantum information from decoherence, noise, and gate imperfections. Extracting the underlying logical state from these codes requires dedicated decoders that infer the likely errors from the measured syndromes and prescribe corrections (dennis2002topological). However, ML decoding for quantum codes is NP-hard (kuo2020hardnesses), prompting the adoption of approximate methods that trade optimality for computational tractability (demarti2024decoding). Classical quantum decoding approaches include MWPM, which achieves near-optimal thresholds under independent noise but suffers from poor scaling even with practical approximations (edmonds1965paths; fowler2012towards; meinerz2022scalable); belief propagation, effective for sparse parity-check codes but impeded by quantum degeneracy (pearl2022fusion; panteleev2021quantum; wang2024artificial); union-find decoders that map syndromes to graph problems but achieve lower thresholds than MWPM (delfosse2021almost); and tensor-network decoders that attain the highest accuracy at steep computational cost (bravyi2014efficient; google2023suppressing). Despite their foundational role, these conventional approaches exhibit inherent limitations that impede practical deployment in large-scale, fault-tolerant quantum systems (krenn2023artificial; demarti2024decoding). Machine learning has emerged as a compelling alternative, with various architectures that demonstrate accuracy and speed gains over classical baselines while allowing adaptation to device-specific, correlated noise processes that challenge traditional decoders (wang2024artificial; demarti2024decoding; varsamopoulos2017decoding; varsamopoulos2019comparing; harper2020efficient; magesan2020effective; liu2019neural). Specifically, these architectures are employed in reinforcement learning, (colomer2020reinforcement; sweke2020reinforcement; fitzek2020deep; ccelikkanat2022piecewise; veeresh2024rl), and supervised learning (bishop2006pattern; goodfellow2016deep), where models are trained on labeled datasets to map measured syndromes to recovery operations (demarti2024decoding; wang2024artificial). Early approaches included feedforward networks (torlai2017neural), neural decoders learning error distributions (krastanov2017deep), and quantum autoencoders (locher2023quantum), demonstrating generalization while reducing complexity and adapting to noise. CNN-based decoders achieve strong performance on topological codes via spatial correlations (maskara2019advantages; meinerz2022scalable). More recently, transformer-based architectures have been explored, most notably the QECCT (choukroun2024deep), outperforming MWPM across topological codes. Another innovative AI-based decoder is AlphaQubit (bausch2024learning), which represents a major milestone in QEC decoding but employs a recurrent structure and processes analog measurement data, unlike our feed-forward architecture which utilizes discrete binary syndrome inputs.
@@ 3 Background
Stabilizer Formalism. The stabilizer framework, (gottesman1997stabilizer), addresses QEC challenges by discretizing the error space while preserving quantum coherence. This formalism exploits the algebraic structure of the Pauli group to construct quantum codes syndrome extraction. The Pauli group foundation. The $n$-qubit Pauli group captures all local quantum errors:
@@ 4 SAQ Decoder
The dual-stream representation construction stage takes syndrome measurements as input and generates initial logical class estimates. These estimates, along with the original syndrome measurements, are then transformed into two token streams that serve as input to the SLTD. Using shared transformer weights, the SLTD processes these streams with distinct attention patterns tailored for QEC: syndrome tokens capture local correlations between neighboring stabilizer measurements, while logical tokens integrate information globally to determine error classes. The dual-stream architecture explicitly models the asymmetric information flow in quantum decoding, from local syndrome violations to global logical error determination. This dual-stream approach, shares high-level similarity with architectures used for classical codes, such as CrossMPT (park2024crossmpt). The SLTD outputs logical class predictions and qubit flip predictions, which are trained using differentiable logical centric losses that approximate discrete GF(2), before being fed to the CPND stage. Subsequently, the CPND enforces syndrome consistency while preserving the transformer’s learned representations, ensuring valid QEC.
@@
where $4^{k}$ represents the total number of logical equivalence classes for $k$ logical qubits. The shallow MLP $b_{\phi}(\mathbf{s})$ provides informed priors about the most likely logical class, enabling the SLTD to refine these estimates rather than exploring the entire logical space from scratch. Such a mapping, where a syndrome input is processed by a shallow feed-forward network, appeared in earlier works, notably the FFN layer in (meinerz2022scalable) and the initial noise estimator in QECCT (choukroun2024deep). Crucially, in our work, $b_{\phi}$ performs a global estimation of the logical class ($\tilde{\bm{\ell}}$), serving as a global prior input to the Logical Stream ($\mathbf{T}_{L}$). This contrasts with related approaches that utilize these initial layers primarily for generating local physical error probabilities or extracting an initial prediction of the recovery operator. The design choice aligns with the stabilizer formalism, where error correction decisions are made purely based on syndrome information, independent of the protected quantum information. With both syndrome measurements $\mathbf{s}$ and initial logical class estimates $\tilde{\bm{\ell}}$ available, we construct two complementary token streams that capture different aspects of the quantum decoding problem:
@@ 4.2 Stage 2: Syndrome-Logical Transformer Decoder (SLTD)
Having constructed dual token streams, $\mathbf{T}_{S}^{[0]}$ and $\mathbf{T}_{L}^{[0]}$, the SLTD refines these representations through $N$ transformer layers with asymmetric attention. We detail each computational step for an arbitrary layer $l$. Both token streams first undergo layer normalization (ba2016layer) before attention computation. The normalized tokens are processed using an asymmetric attention mechanism that captures the fundamental information flow in QEC: syndrome measurements reflect local physical constraints, while logical error determination requires global integration. This design restricts syndrome attention to topological neighborhoods while allowing logical tokens global access, enabling efficient local-global information processing. Syndrome self-attention processes syndrome tokens through:
@@
Logical tokens employ unrestricted attention patterns, enabling global syndrome integration. Following attention computation, residual connections combine the attention outputs with input tokens. Both streams pass through standard FFNs with $4\times$ expansion and GELU activation (hendrycks2016gaussian). Finally, a second residual connection yields the layer outputs. This process transforms initial token representations into refined syndrome and logical embeddings that capture both local correlations and global quantum code structure.
@@ 5 Experiments and Results
Our method was evaluated on various stabilizer codes and noise types, demonstrating the generality of the SAQ-Decoder framework. We primarily focused on topological codes due to their prominence in fault-tolerant quantum computing, specifically toric codes (kitaev1997quantum) and rotated surface codes (bombin2007optimal). To further demonstrate the framework’s broad applicability, we included an evaluation of the Repetition Code and the Color Code using stim (gidney2021stim). We evaluated our method under three well-studied noise models: independent noise, depolarizing noise and circuit noise. Detailed descriptions of the code constructions and stabilizer geometries are provided in Appendix C, training details and hyperparameters are provided in Reproducibility Statement. We evaluate our approach against three key baselines: the QECCT (choukroun2024deep), a state-of-the-art neural decoder that outperforms classical methods, Belief Propagation with Order-2 Ordered Statistics Decoder (BPOSD-2) (roffe2020decoding), and MWPM algorithm (fowler2013minimum), the gold standard classical decoder for surface codes. The BPOSD family of decoders is widely used, although its worst-case complexity scales as $\mathcal{O}(n^{3})$(demarti2024decoding), our implementation leverages the optimized implementation (Roffe_LDPC_Python_tools_2022) to provide a strong, practical classical benchmark for quantum codes. Although MWPM has a worst-case complexity $\mathcal{O}(n^{3}\log n)$ (demarti2024decoding), we use the optimized implementation from higgott2022pymatching which achieves near-quadratic complexity and serves as the primary classical benchmark for topological codes. We also consider the performance of the raw Syndrome Stream prediction, termed SAQ-Decoder (No CPND), as an architectural baseline. As our primary evaluation metric, we use the LER, which measures the probability that the QEC process fails to properly recover the encoded logical information. We also evaluate the code threshold, i.e., the critical noise rate below which increasing code distance improves performance. Together, these metrics assess decoder effectiveness across various noise models, characterizing both absolute performance and scalability.
@@ 5.1 Experimental results
The experiments span code lengths from $L_{\text{code}}=3$ to $11$, comparable to those evaluated in QECCT. These results demonstrate that SAQ-Decoder achieves superior decoding performance compared to state-of-the-art baselines across diverse QEC scenarios. Figure 2 presents the performance of toric codes under depolarizing noise, where SAQ-Decoder exhibits striking advantages. Our method demonstrates consistent superiority across all code distances, with particularly dramatic improvements at $L_{\text{code}}=10$ where SAQ-Decoder achieves $25-50\%$ lower LER compared to MWPM and and BPOSD-2 at physical error rates above 0.15. The scalability benefits are clearly evident as the performance gap widens from $L_{\text{code}}=6$ to $L_{\text{code}}=10$. Similarly, under independent noise in Figure 3, SAQ-Decoder maintains robust performance across all code sizes, with particularly notable advantages at $L_{\text{code}}=8$ and $L_{\text{code}}=10$. Conversely, QECCT exhibits minimal performance gains relative to MWPM and BPOSD-2. Figure 4 demonstrates our method’s performance on rotated surface codes under independent and depolarizing noise models. Under depolarization noise in Figures 4(a)4(b), SAQ-Decoder consistently maintains lower LER than QECCT, BPOSD-2 and MWPM across the entire noise range. Under independent noise conditions in Figures 4(c)4(d), SAQ-Decoder demonstrates even more substantial improvements, while QECCT exhibits inferior performance compared to MWPM and BPOSD-2. These results validate the effectiveness of our learned decoder with post-processing approach across the spectrum of topological QEC codes and noise models. We attribute SAQ-Decoder’s superior performance over QECCT to two key factors: (i) while QECCT focuses on reducing Bit Error Rate (BER), which is not the primary objective in QEC, SAQ-Decoder directly optimizes for logical error suppression; and (ii) our novel architecture combines direct logical class prediction with qubit flip priors, integrating these observations in a post-processing stage that guarantees syndrome consistency—unlike QECCT, which only predicts qubit-level flips without ensuring this crucial constraint. Crucially, the SAQ-Decoder (No CPND) variant consistently outperforms all classical and neural baselines across the entire noise spectrum, demonstrating the power of the dual-stream architecture alone. The SAQ-Decoder achieves error thresholds of 10.99% and 10.7% for toric and rotated surface codes respectively under independent noise, and error thresholds of 18.6% and 18.3% under depolarizing noises. For the toric code under depolarizing noise (Figure 5(a)), this threshold of 18.6% approaches the ML bound of 18.9% (bombin2012strong) while maintaining linear complexity in syndrome size. This significantly outperforms BPOSD-2 and MWPM (16%) (wang2009threshold) and exceeds the previous QECCT result (17.8%). For toric codes under independent noise (Figure 5(b)), we achieve a threshold of 10.99% while significantly outperforming BPOSD-2 (10.8%) and MWPM (10.3%) (wang2003confinement; higgott2022pymatching), essentially reaching the ML threshold estimated between 10.9% and 11.0% (eczoo_toric). For rotated surface codes, SAQ-Decoder demonstrates remarkable consistency with the toric code performance. Under depolarizing noise, we achieve a threshold of 18.3% (Figure 5(c)), significantly exceeding both QECCT (17.2%), BPOSD-2 (14.1%, based on our experiments) and MWPM (14.0%, (demarti2024decoding)), while under independent noise, the threshold reaches 10.7% (Figure 5(d)) where QECCT achieved 10.3%, BPOSD-2 10.2% and MWPM 10.6%, based on our experiments. These findings indicate that our approach generalizes effectively across different topological code geometries without compromising performance. A detailed comparison of our depolarizing noise threshold against other neural and classical decoders is provided in Appendix E.

### v2

@@ 1 Introduction
Since Feynman’s 1982 vision of quantum computation Feynman (2018), significant progress has demonstrated that quantum computers can leverage quantum mechanical principles to achieve fundamental computational advantages over classical methods (Steane, 1998; Ladd et al., 2010; Preskill, 2012; deMarti iOlius et al., 2024). Landmark quantum algorithms have demonstrated computational advantages, including exponential speedup for factoring (Shor, 1994) and quadratic search improvement (Grover, 1996). Recent experimental demonstrations of quantum supremacy have further validated quantum computing’s potential across diverse domains (Arute et al., 2019; Zhong et al., 2020; Wu et al., 2021; Huang et al., 2022; Madsen et al., 2022; Bao et al., 2023; Bluvstein et al., 2024; Aghaee Rad et al., 2025). These advances promise to revolutionize cryptography (Ekert, 1991; Bennett and Brassard, 2014), optimization (Kadowaki and Nishimori, 1998; Bharti et al., 2022), materials science (Lloyd, 1996), and machine learning (Huang et al., 2022; Cerezo et al., 2022).
@@
Yet, for practical quantum computation to become a reality, errors on the physical level must be corrected with high confidence. Despite recent advances, quantum noise remains a major obstacle to practical quantum computing (73). These errors arise through numerous mechanisms: quantum gates cause unwanted errors due to imprecise implementation (Fowler et al., 2012a), while additional errors stem from imperfections in the equipment (Preskill, 2018), interaction with the surrounding environment (Burnett et al., 2019; Etxezarreta Martinez et al., 2021), or measuring quantum systems (Fowler et al., 2012a). While fault-tolerant quantum computation can theoretically be achieved through redundancy by combining multiple physical qubits into one logical qubit (Shor, 1995; Nielsen and Chuang, 2010), this approach creates a critical computational bottleneck: QEC requires real-time decoding algorithms that must process syndrome measurements and determine corrections within microsecond timescales while maintaining near-optimal accuracy (Terhal, 2015; Higgott, 2022). Current decoding methods face a fundamental trade-off between computational efficiency and error-correction performance, methods like MWPM (Fowler, 2013), Belief Propagation with Ordered Statistics Decoding (BP-OSD) (Roffe et al., 2020) and tensor network decoder (Bravyi et al., 2014) scaling prohibitively with code distance while faster heuristics sacrifice the accuracy essential for fault-tolerant operation (deMarti iOlius et al., 2024). The field of QEC has advanced significantly, with several families of QEC codes proposed, including topological codes (Kitaev, 2003; Bombin and Martin-Delgado, 2006; Fowler et al., 2012a; Chamberland et al., 2020), Quantum Low-Density Parity Check (QLDPC) codes (MacKay et al., 2004; Panteleev and Kalachev, 2021; Breuckmann and Eberhardt, 2021), and quantum turbo codes (Poulin et al., 2009). Recently, there has been significant growth in machine learning techniques applied to quantum decoding (Wang and Tang, 2024; Klusch et al., 2024). However, existing neural decoders typically fail to achieve near-optimal error thresholds, creating a gap between the theoretical potential of learned approaches and the performance requirements of fault-tolerant quantum computing. We address this challenge by introducing a unified framework that combines transformer-based neural decoding with specialized architectural innovations. Our approach leverages neural networks to learn syndrome-to-error map- pings while employing dual-stream processing and logical-centric loss design to directly optimize logical error suppression. To achieve this, our framework introduces several key innovations:
@@
  • A novel dual-stream transformer architecture (Vaswani et al., 2017) that simultaneously processes syndrome and logical information streams with specialized attention mechanisms, featuring global tokens (Zaheer et al., 2020) and structured masking patterns that capture the geometric constraints and local correlations inherent in stabilizer codes.
@@
Our results significantly outperform existing neural decoders like QEC Transformer (QECCT) (Choukroun and Wolf, 2024) and classical methods like MWPM, while matching the performance of computationally expensive approaches across both toric and rotated surface codes.
@@ 2 Related works
A broad suite of QEC codes has been devised to protect quantum information from decoherence, noise, and gate imperfections. Extracting the underlying logical state from these codes requires dedicated decoders that infer the likely errors from the measured syndromes and prescribe corrections (Dennis et al., 2002). However, ML decoding for quantum codes is NP-hard (Kuo and Lu, 2020), prompting the adoption of approximate methods that trade optimality for computational tractability (deMarti iOlius et al., 2024). Classical quantum decoding approaches include MWPM, which achieves near-optimal thresholds under independent noise but suffers from poor scaling even with practical approximations (Edmonds, 1965; Fowler et al., 2012b; Meinerz et al., 2022); belief propagation, effective for sparse parity-check codes but impeded by quantum degeneracy (Pearl, 2022; Panteleev and Kalachev, 2021; Wang and Tang, 2024); union-find decoders that map syndromes to graph problems but achieve lower thresholds than MWPM (Delfosse and Nickerson, 2021); and tensor-network decoders that attain the highest accuracy at steep computational cost (S. Bravyi, M. Suchara, and A. Vargo (2014); 73). Despite their foundational role, these conventional approaches exhibit inherent limitations that impede practical deployment in large-scale, fault-tolerant quantum systems (Krenn et al., 2023; deMarti iOlius et al., 2024). Machine learning has emerged as a compelling alternative, with various architectures that demonstrate accuracy and speed gains over classical baselines while allowing adaptation to device-specific, correlated noise processes that challenge traditional decoders (Wang and Tang, 2024; deMarti iOlius et al., 2024; Varsamopoulos et al., 2017; 2019; Harper et al., 2020; Magesan and Gambetta, 2020; Liu and Poulin, 2019). Specifically, these architectures are employed in reinforcement learning, (Colomer et al., 2020; Sweke et al., 2020; Fitzek et al., 2020; Çelikkanat et al., 2022; Veeresh et al., 2024), and supervised learning (Bishop and Nasrabadi, 2006; Goodfellow et al., 2016), where models are trained on labeled datasets to map measured syndromes to recovery operations (deMarti iOlius et al., 2024; Wang and Tang, 2024). Early approaches included feedforward networks (Torlai and Melko, 2017), neural decoders learning error distributions (Krastanov and Jiang, 2017), and quantum autoencoders (Locher et al., 2023), demonstrating generalization while reducing complexity and adapting to noise. CNN-based decoders achieve strong performance on topological codes via spatial correlations (Maskara et al., 2019; Meinerz et al., 2022). More recently, transformer-based architectures have been explored, most notably the QECCT (Choukroun and Wolf, 2024), outperforming MWPM across topological codes. Another innovative AI-based decoder is AlphaQubit (Bausch et al., 2024), which represents a major milestone in QEC decoding but employs a recurrent structure and processes analog measurement data, unlike our feed-forward architecture which utilizes discrete binary syndrome inputs.
@@ 3 Background
Stabilizer Formalism. The stabilizer framework, (Gottesman, 1997), addresses QEC challenges by discretizing the error space while preserving quantum coherence. This formalism exploits the algebraic structure of the Pauli group to construct quantum codes syndrome extraction. The Pauli group foundation. The $n$-qubit Pauli group captures all local quantum errors:
@@ 4 SAQ Decoder
The dual-stream representation construction stage takes syndrome measurements as input and generates initial logical class estimates. These estimates, along with the original syndrome measurements, are then transformed into two token streams that serve as input to the SLTD. Using shared transformer weights, the SLTD processes these streams with distinct attention patterns tailored for QEC: syndrome tokens capture local correlations between neighboring stabilizer measurements, while logical tokens integrate information globally to determine error classes. The dual-stream architecture explicitly models the asymmetric information flow in quantum decoding, from local syndrome violations to global logical error determination. This dual-stream approach, shares high-level similarity with architectures used for classical codes, such as CrossMPT (Park et al., 2024). The SLTD outputs logical class predictions and qubit flip predictions, which are trained using differentiable logical centric losses that approximate discrete GF(2), before being fed to the CPND stage. Subsequently, the CPND enforces syndrome consistency while preserving the transformer’s learned representations, ensuring valid QEC.
@@
where $4^{k}$ represents the total number of logical equivalence classes for $k$ logical qubits. The shallow MLP $b_{\phi}(\mathbf{s})$ provides informed priors about the most likely logical class, enabling the SLTD to refine these estimates rather than exploring the entire logical space from scratch. Such a mapping, where a syndrome input is processed by a shallow feed-forward network, appeared in earlier works, notably the FFN layer in (Meinerz et al., 2022) and the initial noise estimator in QECCT (Choukroun and Wolf, 2024). Crucially, in our work, $b_{\phi}$ performs a global estimation of the logical class ($\tilde{\bm{\ell}}$), serving as a global prior input to the Logical Stream ($\mathbf{T}_{L}$). This contrasts with related approaches that utilize these initial layers primarily for generating local physical error probabilities or extracting an initial prediction of the recovery operator. The design choice aligns with the stabilizer formalism, where error correction decisions are made purely based on syndrome information, independent of the protected quantum information. With both syndrome measurements $\mathbf{s}$ and initial logical class estimates $\tilde{\bm{\ell}}$ available, we construct two complementary token streams that capture different aspects of the quantum decoding problem:
@@ 4.2 Stage 2: Syndrome-Logical Transformer Decoder (SLTD)
Having constructed dual token streams, $\mathbf{T}_{S}^{[0]}$ and $\mathbf{T}_{L}^{[0]}$, the SLTD refines these representations through $N$ transformer layers with asymmetric attention. We detail each computational step for an arbitrary layer $l$. Both token streams first undergo layer normalization (Ba et al., 2016) before attention computation. The normalized tokens are processed using an asymmetric attention mechanism that captures the fundamental information flow in QEC: syndrome measurements reflect local physical constraints, while logical error determination requires global integration. This design restricts syndrome attention to topological neighborhoods while allowing logical tokens global access, enabling efficient local-global information processing. Syndrome self-attention processes syndrome tokens through:
@@
Logical tokens employ unrestricted attention patterns, enabling global syndrome integration. Following attention computation, residual connections combine the attention outputs with input tokens. Both streams pass through standard FFNs with $4\times$ expansion and GELU activation (Hendrycks and Gimpel, 2016). Finally, a second residual connection yields the layer outputs. This process transforms initial token representations into refined syndrome and logical embeddings that capture both local correlations and global quantum code structure.
@@ 5 Experiments and Results
Our method was evaluated on various stabilizer codes and noise types, demonstrating the generality of the SAQ-Decoder framework. We primarily focused on topological codes due to their prominence in fault-tolerant quantum computing, specifically toric codes (Kitaev, 1997) and rotated surface codes (Bombín and Martin-Delgado, 2007). To further demonstrate the framework’s broad applicability, we included an evaluation of the Repetition Code and the Color Code using stim (Gidney, 2021). We evaluated our method under three well-studied noise models: independent noise, depolarizing noise and circuit noise. Detailed descriptions of the code constructions and stabilizer geometries are provided in Appendix C, training details and hyperparameters are provided in Reproducibility Statement. We evaluate our approach against three key baselines: the QECCT (Choukroun and Wolf, 2024), a state-of-the-art neural decoder that outperforms classical methods, Belief Propagation with Order-2 Ordered Statistics Decoder (BPOSD-2) (Roffe et al., 2020), and MWPM algorithm (Fowler, 2013), the gold standard classical decoder for surface codes. The BPOSD family of decoders is widely used, although its worst-case complexity scales as $\mathcal{O}(n^{3})$(deMarti iOlius et al., 2024), our implementation leverages the optimized implementation (Roffe, 2022) to provide a strong, practical classical benchmark for quantum codes. Although MWPM has a worst-case complexity $\mathcal{O}(n^{3}\log n)$ (deMarti iOlius et al., 2024), we use the optimized implementation from Higgott (2022) which achieves near-quadratic complexity and serves as the primary classical benchmark for topological codes. We also consider the performance of the raw Syndrome Stream prediction, termed SAQ-Decoder (No CPND), as an architectural baseline. As our primary evaluation metric, we use the LER, which measures the probability that the QEC process fails to properly recover the encoded logical information. We also evaluate the code threshold, i.e., the critical noise rate below which increasing code distance improves performance.
@@ 5.1 Experimental results
The experiments span code lengths from $L_{\text{code}}=3$ to $11$, comparable to those evaluated in QECCT. These results demonstrate that SAQ-Decoder achieves superior decoding performance compared to state-of-the-art baselines across diverse QEC scenarios. Figure 2 presents the performance of toric codes under depolarizing noise, where SAQ-Decoder exhibits striking advantages. Our method demonstrates consistent superiority across all code distances, with particularly dramatic improvements at $L_{\text{code}}=10$ where SAQ-Decoder achieves $25-50\%$ lower LER compared to MWPM and and BPOSD-2 at physical error rates above 0.15. The scalability benefits are clearly evident as the performance gap widens from $L_{\text{code}}=6$ to $L_{\text{code}}=10$. Similarly, under independent noise in Figure 3, SAQ-Decoder maintains robust performance across all code sizes, with particularly notable advantages at $L_{\text{code}}=8$ and $L_{\text{code}}=10$. Conversely, QECCT exhibits minimal performance gains relative to MWPM and BPOSD-2. Figure 4 demonstrates our method’s performance on rotated surface codes under independent and depolarizing noise models. Under depolarization noise in Figures 4(a)4(b), SAQ-Decoder consistently maintains lower LER than QECCT, BPOSD-2 and MWPM across the entire noise range. Under independent noise conditions in Figures 4(c)4(d), SAQ-Decoder demonstrates even more substantial improvements, while QECCT exhibits inferior performance compared to MWPM and BPOSD-2. These results validate the effectiveness of our learned decoder with post-processing approach across the spectrum of topological QEC codes and noise models. We attribute SAQ-Decoder’s superior performance over QECCT to two key factors: (i) while QECCT focuses on reducing Bit Error Rate (BER), which is not the primary objective in QEC, SAQ-Decoder directly optimizes for logical error suppression; and (ii) our novel architecture combines direct logical class prediction with qubit flip priors, integrating these observations in a post-processing stage that guarantees syndrome consistency—unlike QECCT, which only predicts qubit-level flips without ensuring this crucial constraint. Crucially, the SAQ-Decoder (No CPND) variant consistently outperforms all classical and neural baselines across the entire noise spectrum, demonstrating the power of the dual-stream architecture alone. The SAQ-Decoder achieves error thresholds of 10.99% and 10.7% for toric and rotated surface codes respectively under independent noise, and error thresholds of 18.6% and 18.3% under depolarizing noises. For the toric code under depolarizing noise (Figure 5(a)), this threshold of 18.6% approaches the ML bound of 18.9% (Bombin et al., 2012) while maintaining linear complexity in syndrome size. This significantly outperforms BPOSD-2 and MWPM (16%) (Wang et al., 2009) and exceeds the previous QECCT result (17.8%). For toric codes under independent noise (Figure 5(b)), we achieve a threshold of 10.99% while significantly outperforming BPOSD-2 (10.8%) and MWPM (10.3%) (Wang et al., 2003; Higgott, 2022), essentially reaching the ML threshold estimated between 10.9% and 11.0% (76). For rotated surface codes, SAQ-Decoder demonstrates remarkable consistency with the toric code performance. Under depolarizing noise, we achieve a threshold of 18.3% (Figure 5(c)), significantly exceeding both QECCT (17.2%), BPOSD-2 (14.1%, based on our experiments) and MWPM (14.0%, (deMarti iOlius et al., 2024)), while under independent noise, the threshold reaches 10.7% (Figure 5(d)) where QECCT achieved 10.3%, BPOSD-2 10.2% and MWPM 10.6%, based on our experiments. These findings indicate that our approach generalizes effectively across different topological code geometries without compromising performance. A detailed comparison of our depolarizing noise threshold against other neural and classical decoders is provided in Appendix E.